Q)HI was heated in a closed tube at 440 degrees C
till equilibrium is reached . at this temperature 22% of HI was dissociated . The equilibrium constant for this dissociation is -
a)0.282 b)0.0796 c)0.0199 d) 1.99.
 

Grade:12

2 Answers

Priya
3 years ago
Given reaction :
  • 2HI is in equilibrium with H2 and I2
  •                    2HI     H2  I2
​Initial moles:.          2.        -.      -
At equilibrium.        2-2x.  X.     X
X=0.22 ( amount dissociated of HI)
 
So,
K= (x)(x)/(2-2x)^2
K=(0.22)^2/(1.56)^2
K=0.0199 
Arvind
3 years ago
 
Given reaction :
  • 2HI is in equilibrium with H2 and I2
  •                    2HI     H2  I2
​Initial moles:.          2.        -.      -
At equilibrium.        2-2x.  X.     X
X=0.22 ( amount dissociated of HI)
 
So,
K= (x)(x)/(2-2x)^2
K=(0.22)^2/(1.56)^2
K=0.0199  
Given reaction :
  • 2HI is in equilibrium with H2 and I2
  •                    2HI     H2  I2
​Initial moles:.          2.        -.      -
At equilibrium.        2-2x.  X.     X
X=0.22 ( amount dissociated of HI)
 
So,
K= (x)(x)/(2-2x)^2
K=(0.22)^2/(1.56)^2
K=0.0199  
Given reaction :
  • 2HI is in equilibrium with H2 and I2
  •                    2HI     H2  I2
​Initial moles:.          2.        -.      -
At equilibrium.        2-2x.  X.     X
X=0.22 ( amount dissociated of HI)
 
So,
K= (x)(x)/(2-2x)^2
K=(0.22)^2/(1.56)^2
K=0.0199